Two numbers have a sum = 18 and a product = 80. Find the two numbers. List the smaller number first
Answer by TimothyLamb(4379)

x = smaller number
y = larger number
---
x + y = 18
y = 18 - x
xy = 80
x(18 - x) = 80
-xx + 18x - 80 = 0
xx - 18x + 80 = 0
---
the above quadratic equation is in standard form, with a=1, b=-18 and c=80
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -18 80
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = 10
x = 8
---
x = 8
y = 18 - x
y = 18 - 8
y = 10
---
answer:
x = 8
y = 10
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---